OpenStudy (anonymous):

finding unknown side length of triangle problem below vvvv and also telling whether it is a pythagorean triple.

OpenStudy (anonymous):



OpenStudy (anonymous):

the 8 is actually 8(square root)2

OpenStudy (anonymous):

Try applying the Pythagorean theorem to it (a^2+b^2=c^2), where a and b are the lengths of the two sides, and c is the length of the hypotenuse. In this case, a and b are equal therefore you can change the equation to 2a^2=c^2 . Substitute in the c-value, and rearrange to solve for a (the x-value in your diagram).

OpenStudy (anonymous):

all im given is that the one side is \[8\sqrt{2}\]

OpenStudy (nincompoop):

could it be that your two sides are sort(32)

OpenStudy (nincompoop):

which means that you have sqrt(4^2 * 2)

OpenStudy (nincompoop):

stupid auto-correct put it as sort LOL

OpenStudy (anonymous):

You can substitute \[8\sqrt{3} \]into the c-value. \[2a^2=(8\sqrt{3})^2\] If you sqrt the right side, you'll get \[2a^2=(8\sqrt{3})^2\] = 64(3) = 192 Then rearrange to solve for a.

OpenStudy (anonymous):

its 2 squared but thanks i got it

OpenStudy (anonymous):

or square root 2*